# Wasserstein Evolution : Evolutionary Optimization as Phase Transition

### v1

@@ 4  Experiments
To empirically validate the theoretical framework and algorithmic performance of Wasserstein Evolution (WE), we conduct comprehensive experiments on five classical benchmark functions: Rastrigin, Beale, Himmelblau, Six-Hump Camel Back, and Holder Table. These functions represent diverse optimization challenges, including high-dimensional multimodality, symmetry, and deceptive landscapes. We compare WE against three well-established evolutionary algorithms: Genetic Algorithm (GA), Differential Evolution (DE), and Covariance Matrix Adaptation Evolution Strategy (CMA-ES).
@@ 4.1  Experimental Setup
We conducted experiments on five classical benchmark functions to evaluate the Wasserstein Evolution (WE) against three established evolutionary algorithms: Genetic Algorithm (GA), Differential Evolution (DE), and Covariance Matrix Adaptation Evolution Strategy (CMA-ES). These functions, detailed in Table 1, were selected for their diverse characteristics, including multimodality, symmetry, and deceptive landscapes, which provide a comprehensive testbed for assessing exploration-exploitation balance.
Table 1: Benchmark Functions Used in the Experiments
Name Formula Features
Rastrigin $f(\mathbf{x})=10n+\sum_{i=1}^{n}\left[x_{i}^{2}-10\cos(2\pi x_{i})\right]$, $\mathbf{x}\in[-5.12,5.12]^{n}$ ($n=2$). Highly multimodal, separable. Global optimum: $f(\mathbf{0})=0$.
Beale $f(x,y)=(1.5-x+xy)^{2}+(2.25-x+xy^{2})^{2}+(2.625-x+xy^{3})^{2}$, $x,y\in[-4.5,4.5]$. Non-convex, unimodal with a narrow valley. Global optimum: $f(3,0.5)=0$.
Himmelblau $f(x,y)=(x^{2}+y-11)^{2}+(x+y^{2}-7)^{2}$, $x,y\in[-6,6]$. Symmetric with four equal-valued global minima. Minima: $(3.0,2.0)$, (2.805,3.131)(-2.805,3.131), (3.779,3.283)(-3.779,-3.283), (3.584,1.848)(3.584,-1.848), $f^{*}=0$.
Six-Hump Camel Back $f(x,y)=\left(4-2.1x^{2}+\frac{x^{4}}{3}\right)x^{2}+xy+(-4+4y^{2})y^{2}$, $x\in[-3,3],y\in[-2,2]$. Two global minima and four local minima. Global minima: $f(0.0898,-0.7126)=f(-0.0898,0.7126)\approx-1.0316$.
Holder Table $f(x,y)=-\left|\sin(x)\cos(y)\exp\left(\left|1-\frac{\sqrt{x^{2}+y^{2}}}{\pi}\right|\right)\right|$, $x,y\in[-10,10]$. Highly multimodal with numerous local minima. Four global minima: $f(\pm 8.05502,\pm 9.66459)\approx-19.2085$.
All algorithms were configured with a population size of $N=30$ and executed for $K=500$ iterations. For the proposed WE algorithm, we set the initial inverse temperature $\beta_{0}=0.1$, the final inverse temperature $\beta_{\text{final}}=1000$, the base learning rate $\eta_{0}=0.1$ with a linear decay schedule, and the kernel density estimation (KDE) bandwidth $h=0.5$.
Performance is assessed using two complementary metrics. The first is the Potential Energy, defined as the best function value $f(\mathbf{x}^{*})$ found by the algorithm, which quantifies its exploitation success. The second is the Differential Entropy, which quantifies the population diversity. For a final population $\{\mathbf{x}_{i}\}_{i=1}^{N}$, the entropy $S$ is estimated via multivariate Gaussian Kernel Density Estimation (KDE):
@@
where $\hat{\rho}$ is the KDE estimate. A higher entropy value indicates a more diverse, multi-modal population distribution, reflecting stronger exploration capacity.
4.2  Results and Analysis
@@
The performance of all four algorithms across the five benchmark functions is summarized in Table 2. Each cell reports the final potential energy $U$ (corresponding to the optimal function value) and the corresponding differential entropy $S$.
Table 2: WE and Its Competitors’ Performance on Potential Energy $U$ and Entropy $S$
Function WE GA DE CMA-ES
Rastrigin $U$: 0.00, $S$: 3.57 $U$: 0.00, $S$: -34.15 $U$: 0.00, $S$: -34.26 $U$: 0.00, $S$: -34.63
Beale $U$: 0.00, $S$: -3.19 $U$: 0.02, $S$: -21.44 $U$: 0.00, $S$: -33.92 $U$: 0.00, $S$: -33.43
Himmelblau $U$: 0.00, $S$: 3.98 $U$: 0.06, $S$: -34.08 $U$: 0.00, $S$: -34.05 $U$: 0.00, $S$: -2.16
Six-Hump Camel Back $U$: -1.03, $S$: 1.54 $U$: -1.03, $S$: -34.23 $U$: -1.03, $S$: -34.23 $U$: -1.03, $S$: -3.38
Holder Table $U$: -19.21, $S$: 6.19 $U$: -19.21, $S$: -34.86 $U$: -19.14, $S$: -34.02 $U$: -19.18, $S$: 5.98
Fig. 5: Final Population Distribution of Different Algorithms
The results reveal several key insights:
1. Superior Entropy Preservation: WE consistently achieves differential entropy values that are orders of magnitude higher than GA and DE, and significantly higher than CMA-ES on most functions. This empirically validates WE’s core mechanism of maintaining exploration pressure through the entropic force term $\frac{1}{\beta}\nabla\log\rho_{t}(\mathbf{x})$.
2. Competitive Convergence: Despite maintaining high entropy (diversity), WE matches or outperforms the other algorithms in finding the global minimum (or a very good one) across all tested functions. This demonstrates that the Wasserstein gradient flow of free energy $\mathcal{F}_{\beta}[\rho]$ successfully navigates the trade-off between exploration and exploitation.
3. Theoretical Alignment with Boltzmann Distribution: The observed behavior is a direct consequence of the theoretical steady-state of the Wasserstein gradient flow. Minimizing the free energy functional $\mathcal{F}_{\beta}[\rho]=\mathbb{E}_{\rho}[f]-\frac{1}{\beta}S[\rho]$ leads to the Gibbs-Boltzmann distribution:
$\rho_{\beta}(\mathbf{x})=\frac{1}{Z(\beta)}\exp(-\beta f(\mathbf{x})),\qquad Z(\beta)=\int\exp(-\beta f(\mathbf{x}))d\mathbf{x}.$ (18)
WE’s dynamics drive the population distribution $\rho_{t}$ towards this $\rho_{\beta}$, explaining why it naturally maintains diversity on multi-modal functions. The annealing schedule for $\beta_{k}$ controls the phase transition: at low $\beta$ (high temperature), $\rho_{\beta}$ is diffuse (high entropy, exploration); at high $\beta$ (low temperature), $\rho_{\beta}$ concentrates on minima (low entropy, exploitation). 4. Implications for Multi-modal Optimization: The significantly higher entropy values for WE indicate that its final population is not collapsed to a single point or a small cluster but remains distributed across multiple basins of attraction. This is particularly valuable for real-world optimization problems where identifying multiple good solutions (global or local) is often desired, rather than just a single optimum.
In conclusion, the experiments confirm that Wasserstein Evolution effectively operationalizes the theoretical concept of free energy minimization via gradient flow. It achieves a fundamentally different balance between exploration and exploitation compared to classical evolutionary algorithms, resulting in superior diversity preservation without compromising solution quality, precisely aligning with the behavior predicted by its underlying statistical physics framework.

### v2

@@ 4.1  Experimental Setup
We conducted comprehensive experiments to evaluate Wasserstein Evolution (WE) against five established evolutionary algorithms: Genetic Algorithm (GA)[17], Differential Evolution (DE)[18], Covariance Matrix Adaptation Evolution Strategy (CMA-ES)[19], Adaptive Differential Evolution with Optional External Archive (JADE)[20], and Self-adaptive Differential Evolution (SaDE)[21].
The test suite comprises ten functions, categorized into two types: five classical mathematical benchmark functions and five physical potential energy functions. The benchmark functions were selected for their diverse characteristics including multimodality, symmetry, and deceptive landscapes. The physical potential functions model complex real-world systems such as crystal adsorption, molecular conformation transitions, magnetic confinement plasmas, ion traps, and optical lattices. These two categories provide complementary testbeds for assessing algorithmic performance on both abstract mathematical landscapes and physically meaningful energy surfaces. Details of all test functions are provided in Table 1. All algorithms were configured with a population size of $N=30$ and executed for $K=500$ iterations, with each experiment repeated 30 times for statistical reliability.
Performance is assessed using two complementary metrics computed from the final population. The first is the Average Free Energy $\mathcal{F}$, calculated as:
$\mathcal{F}=\mathbb{E}_{\rho}[U]-\frac{1}{\beta}S$ (27)
where $U(\mathbf{x}_{i})$ is the objective function value (potential energy) and $\beta$ is the final inverse temperature. This composite metric directly reflects the algorithm’s ability to balance energy minimization with entropy preservation.
The second is the Average Differential Entropy $S$, which quantifies the population’s configurational diversity. For a final population $\{\mathbf{x}_{i}\}_{i=1}^{N}$, the entropy is estimated via multivariate Gaussian Kernel Density Estimation (KDE):
@@
where $\hat{\rho}$ is the KDE estimate. Higher entropy values indicate more diverse, multi-modal population distributions, reflecting stronger exploration capacity. Together, these metrics provide a comprehensive evaluation framework grounded in statistical physics principles.
Table 1: Benchmark and Physical Potential Functions Used in the Experiments
Category Name & Formula Domain Features / Physical Interpretation
Benchmark Rastrigin: $f(\mathbf{x})=10n+\sum_{i=1}^{n}\left[x_{i}^{2}-10\cos(2\pi x_{i})\right]$ $5.12,5.12]^{2}$ Highly multimodal, separable. Global optimum: $f(\mathbf{0})=0$.
Beale: $f(x,y)=(1.5-x+xy)^{2}+(2.25-x+xy^{2})^{2}+(2.625-x+xy^{3})^{2}$ $4.5,4.5]^{2}$ Non-convex, unimodal with a narrow valley. Global optimum: $f(3,0.5)=0$.
Himmelblau: $f(x,y)=(x^{2}+y-11)^{2}+(x+y^{2}-7)^{2}$ $6,6]^{2}$ Symmetric with four equal-valued global minima ($f^{*}=0$).
Six-Hump Camel: $f(x,y)=\left(4-2.1x^{2}+\frac{x^{4}}{3}\right)x^{2}+xy+(-4+4y^{2})y^{2}$ $x\in[-3,3],y\in[-2,2]$ Two global minima ($f\approx-1.0316$) and four local minima.
Holder Table: $f(x,y)=-\left|\sin(x)\cos(y)\exp\left(\left|1-\frac{\sqrt{x^{2}+y^{2}}}{\pi}\right|\right)\right|$ $10,10]^{2}$ Highly multimodal. Four global minima ($f\approx-19.2085$).
Physical Potential 2D Periodic: $V=-V_{0}[\cos(\frac{2\pi x}{a})+\cos(\frac{2\pi y}{a})+\lambda_{c}\cos(\frac{2\pi x}{a})\cos(\frac{2\pi y}{a})],V_{0}=2.0,a=1.0,\lambda_{c}=0.5$ $2,2]^{2}$ Periodic potential for crystal adsorption. Symmetric lattice wells.
Double-Well: $V=(x^{2}-1)^{2}+\frac{1}{2}ky^{2}+A_{g}\exp[-B(x^{2}+y^{2})],k=2.0,A_{g}=3.0,B=3.0$ $2,2]^{2}$ Molecular conformation transitions. Two metastable states separated by a Gaussian barrier.
Tokamak: $V=\alpha(r-r_{0})^{2}+\beta(r-r_{0})^{4}+\epsilon(r-r_{0})\cos(m\theta)+\delta\cos(n\theta),r_{0}=2.0,\alpha=0.5,\beta=0.05,\epsilon=0.8,\delta=0.6,m=3,n=2$ $4,4]^{2}$ Charged particle drift in tokamak fields. Magnetic islands create multiple metastable states.
Multipole: $V=\sum_{i=1}^{4}\frac{s_{i}q}{4\pi\epsilon_{0}\sqrt{(x-x_{i})^{2}+(y-y_{i})^{2}+\delta^{2}}},q=1.0,\epsilon_{0}=1.0,\delta=0.3$ $3,3]^{2}$ Quadrupole ion trap potential. Alternating charges create saddle points and trapping regions.
Optical Lattice: $V=V_{0}[\sin^{2}(kx)+\sin^{2}(ky)]\exp[-\frac{x^{2}+y^{2}}{2\sigma^{2}}],V_{0}=3.0,k=2.0,\sigma=3.0$ $4,4]^{2}$ Optical lattice for cold atoms. Periodic modulation with Gaussian envelope creates multiple traps.
4.2  Experiments on Test Functions
@@
The performance of all six algorithms across the ten test functions is summarized in Table 2. Each cell reports the average Free Energy $\mathcal{F}$ and the corresponding average differential Entropy $S$ over 30 independent runs.
Table 2: Average Free Energy $\mathcal{F}$ and Entropy $S$ of WE and Competitors
Function (Metric) WE GA DE CMA-ES JADE SaDE
Rastrigin ($\mathcal{F}$) -0.0039 0.0698 0.0343 0.0342 0.0343 0.0344
Rastrigin ($S$) 3.8783 -34.3251 -34.2552 -34.2164 -34.3120 -34.3626
Beale ($\mathcal{F}$) 0.0018 0.3087 0.0343 0.0343 0.0343 0.0343
Beale ($S$) -1.7029 -33.1249 -34.3191 -34.3006 -34.3421 -34.3226
Himmelblau ($\mathcal{F}$) -0.0039 0.0405 0.0343 0.0090 0.0343 0.0343
Himmelblau ($S$) 3.9380 -34.3633 -34.3011 -5.9218 -34.3114 -34.2870
Six-Hump Camel ($\mathcal{F}$) -1.0332 -0.9971 -0.9973 -1.0271 -0.9979 -0.9973
Six-Hump Camel ($S$) 1.6167 -34.3446 -34.2828 -4.5358 -33.7709 -34.3261
Holder Table ($\mathcal{F}$) -19.2146 -19.1732 -19.0401 -19.1743 -19.1771 -19.1752
Holder Table ($S$) 6.1341 -33.7130 -32.1155 1.2729 -31.3723 -33.2832
2D Periodic ($\mathcal{F}$) -5.0032 -4.9656 -4.9895 -4.9995 -4.9834 -4.9966
2D Periodic ($S$) 3.1528 -34.2458 -10.4928 -0.4719 -16.5898 -3.4461
Double-Well ($\mathcal{F}$) 0.1190 0.1520 0.1502 0.1207 0.1490 0.1520
Double-Well ($S$) -1.2880 -34.3109 -32.5156 -2.9730 -31.2836 -34.3225
Tokamak ($\mathcal{F}$) -0.6927 -0.6559 -0.6570 -0.6898 -0.6570 -0.6563
Tokamak ($S$) 2.8447 -33.9353 -32.8931 -0.0087 -32.8407 -33.4977
Multipole ($\mathcal{F}$) -0.2185 -0.1808 -0.1820 -0.2068 -0.1815 -0.1803
Multipole ($S$) 3.8663 -33.6437 -32.5899 -7.7928 -33.0796 -34.3084
Optical Lattice ($\mathcal{F}$) -0.0042 0.0335 0.0343 -0.0034 0.0317 0.0343
Optical Lattice ($S$) 4.1921 -33.4341 -34.3124 4.0787 -31.7351 -34.3470
The results clearly show that the proposed WE consistently achieves the lowest free energy values across all ten test functions. This aligns with the theoretical intuition that the Wasserstein gradient flow follows the steepest descent path for free energy minimization. Furthermore, WE maintains the highest entropy among all compared algorithms, which is consistent with the theoretical expectation that the optimal population distribution should converge to the Boltzmann distribution at equilibrium. The performance superiority of WE holds for both classical mathematical benchmarks and physically inspired potential functions, demonstrating its robustness and generalizability across diverse and complex landscapes.
Fig. 5: Trajectories of the Population in WE
To further investigate the phase transition behavior of WE and verify its convergence to the Boltzmann distribution, we conducted additional analyses on selected complex landscapes. Taking the Himmelblau, Six-Hump Camel Back, Double-Well, and Tokamak functions as examples, we visualize both the final population distribution and the search trajectories.
Fig. 6: Probability Density Heatmap of the Final Population in WE
Figure 6 presents probability density heatmaps of WE’s final populations.These visualizations confirm that the empirical distribution aligns with the theoretical Boltzmann distribution, as regions with lower energy exhibit higher population density.Notably, the density peaks correspond exactly to the multiple minima of each function, providing empirical evidence that WE effectively samples from the Boltzmann distribution and successfully locates multiple optima in complex landscapes. Figure 5 illustrates the convergence process by showing the search trajectories of individual particles. The trajectories reveal how particles, driven by the combined forces of potential gradient and entropy, gradually navigate from broadly distributed initial positions toward the low-energy regions. This process visually manifests the phase transition from a high-entropy, exploratory state (disorder) to a low-entropy, concentrated state (order) around the optimal solutions, consistent with the theoretical interpretation of optimization as a free-energy-minimization-driven phase transition.
4.3  Translation, Scale, and Rotation Invariance of WE
A robust optimization algorithm should ideally possess certain invariances to basic transformations of the search space. Specifically, translation, scale, and rotation invariance are desirable properties[30, 31]. Translation invariance ensures that an algorithm’s performance does not depend on the absolute location of the optimum in the search space. Scale invariance implies that the algorithm performs consistently regardless of the units or magnitude scaling of the decision variables. Rotation invariance is crucial for handling non-separable problems, as it ensures the algorithm is not biased towards the original coordinate axes. For gradient-based or geometry-aware algorithms like WE, which utilize both the gradient of the objective function ($\nabla f$) and the gradient of the population density ($\nabla\log\rho$), verifying these invariances is essential to confirm that its dynamics are fundamentally tied to the intrinsic geometry of the fitness landscape rather than its specific parameterization.
To systematically test these properties of WE, we employ the Schwefel 2.22 function as a base function and construct a suite of transformed variants, as detailed in Table 3. These variants include pure translation, scaling, rotation, and composite transformations that combine all three operations in different sequences. If WE exhibits the aforementioned invariances, its convergence behavior, final free energy, and entropy should remain statistically consistent across the original function and all its transformed versions, except for predictable shifts or scaling in the solution values themselves. The following experiments and analyses are designed to investigate this hypothesis.
Table 3: Schwefel 2.22 Function and Its Transformations for Invariance Testing
Category Name & Transformation Description Domain Mathematical Form / Notes
Base Function Original Schwefel 2.22 $100,100]^{2}$ $f(\mathbf{x})=\sum_{i=1}^{n}\left(\sum_{j=1}^{i}x_{j}\right)^{2}$, $n=2$. The global minimum is $f(\mathbf{0})=0$.
Translation Schwefel 2.22-Shift Right 20 $100,100]^{2}$ $\mathbf{x}_{\text{trans}}=\mathbf{x}-20$. Minimum at $\mathbf{x}^{*}=(20,20)$.
Schwefel 2.22-Shift Left 30 $100,100]^{2}$ $\mathbf{x}_{\text{trans}}=\mathbf{x}+30$. Minimum at $\mathbf{x}^{*}=(-30,-30)$.
Schwefel 2.22-Shift (15, 15) $100,100]^{2}$ $x_{\text{trans},0}=x_{0}-15$, $x_{\text{trans},1}=x_{1}-15$. Minimum at $\mathbf{x}^{*}=(15,15)$.
Scaling Schwefel 2.22-Scale x2 (Enlarge) $100,100]^{2}$ $\mathbf{x}_{\text{trans}}=\mathbf{x}\times 0.5$. Function landscape is enlarged by a factor of 2.
Schwefel 2.22-Scale x0.5 (Reduce) $100,100]^{2}$ $\mathbf{x}_{\text{trans}}=\mathbf{x}\times 2.0$. Function landscape is shrunk by half.
Schwefel 2.22-Anisotropic Scale $100,100]^{2}$ $x_{1}=x_{1}\times 0.6667$, $x_{2}=x_{2}\times 1.25$. Different scaling per axis.
Rotation Schwefel 2.22-Rotate 45° $100,100]^{2}$ $\mathbf{x}_{\text{trans}}=R(45^{\circ})\,\mathbf{x}$, where $R$ is a 2D rotation matrix.
Schwefel 2.22-Rotate -30° $100,100]^{2}$ $\mathbf{x}_{\text{trans}}=R(-30^{\circ})\,\mathbf{x}$.
Schwefel 2.22-Rotate 75° $100,100]^{2}$ $\mathbf{x}_{\text{trans}}=R(75^{\circ})\,\mathbf{x}$.
Composite Schwefel 2.22-Translate-Rotate-Scale $100,100]^{2}$ Sequence: Translate by (10,10)(10,-10) $\rightarrow$ Rotate 60° $\rightarrow$ Scale $(0.8,1.2)$.
Schwefel 2.22-Rotate-Translate-Scale $100,100]^{2}$ Sequence: Rotate 45° $\rightarrow$ Translate by (15,15)(15,-15) $\rightarrow$ Scale $(0.7,1.5)$.
Fig. 7: Contour plots of the Schwefel 2.22 function and its transformed variants.
Table 4: Average Free Energy $\mathcal{F}$ and Entropy $S$ of WE and Competitors on Schwefel 2.22 Transformations
Function (Metric) WE GA DE CMA-ES JADE SaDE
Original ($\mathcal{F}$) 0.0057 0.4980 0.0343 0.0344 0.0344 0.0343
Original ($S$) -5.7499 -33.7497 -34.3451 -34.3637 -34.3649 -34.3328
Shift Right 20 ($\mathcal{F}$) 0.0057 0.7317 0.0343 0.0344 0.0343 0.0343
Shift Right 20 ($S$) -5.7420 -33.3897 -34.3341 -34.3665 -34.2949 -34.2984
Shift Left 30 ($\mathcal{F}$) 0.0062 0.6788 0.0343 0.0343 0.0343 0.0343
Shift Left 30 ($S$) -6.1653 -33.8295 -34.3042 -34.3201 -34.3101 -34.2928
Shift (15, 15) ($\mathcal{F}$) 0.0058 0.4777 0.0343 0.0344 0.0343 0.0343
Shift (15, 15) ($S$) -5.7807 -33.2944 -34.3173 -34.3684 -34.3241 -34.3355
Scale x2 ($\mathcal{F}$) 0.0076 0.1423 0.0343 0.0343 0.0343 0.0343
Scale x2 ($S$) -7.1719 -33.8689 -34.2644 -34.2716 -34.3068 -34.2746
Scale x0.5 ($\mathcal{F}$) 0.0067 2.0389 0.0343 0.0343 0.0344 0.0344
Scale x0.5 ($S$) -6.6958 -33.7282 -34.2921 -34.3427 -34.3565 -34.3706
Anisotropic Scale ($\mathcal{F}$) 0.0056 0.3047 0.0343 0.0343 0.0344 0.0343
Anisotropic Scale ($S$) -5.6218 -33.7350 -34.3409 -34.2988 -34.3905 -34.3028
Rotate 45° ($\mathcal{F}$) 0.0052 0.2298 0.0343 0.0344 0.0343 0.0344
Rotate 45° ($S$) -5.1908 -33.1372 -34.3493 -34.3539 -34.3287 -34.3611
Rotate -30° ($\mathcal{F}$) 0.0057 0.5717 0.0343 0.0343 0.0343 0.0343
Rotate -30° ($S$) -5.6895 -34.3541 -34.2934 -34.2986 -34.2988 -34.2832
Rotate 75° ($\mathcal{F}$) 0.0060 0.4285 0.0344 0.0343 0.0343 0.0344
Rotate 75° ($S$) -5.9855 -34.3367 -34.3740 -34.3358 -34.3411 -34.3876
Translate-Rotate-Scale ($\mathcal{F}$) 0.0052 0.2658 0.0343 0.0343 0.0344 0.0343
Translate-Rotate-Scale ($S$) -5.1885 -34.3381 -34.3096 -34.3098 -34.3963 -34.3146
Rotate-Translate-Scale ($\mathcal{F}$) 0.0053 0.2223 0.0343 0.0343 0.0343 0.0343
Rotate-Translate-Scale ($S$) -5.3402 -34.3666 -34.3172 -34.3149 -34.2942 -34.2662
The quantitative results presented in Table 4 provide compelling empirical evidence for the translation, scale, and rotation invariance of Wasserstein Evolution (WE). Across all twelve variants of the Schwefel 2.22 function, WE consistently maintains the **lowest free energy** values (all within the narrow range of 0.0052–0.0076) and the highest entropy values (ranging from -5.2 to -7.2). This remarkable stability in both performance metrics, regardless of the applied transformation, directly demonstrates WE’s invariance properties. In stark contrast, while other algorithms like DE, CMA-ES, JADE, and SaDE achieve similarly low free energy values, they do so at the cost of near-zero population diversity, with entropy values consistently around -34.3. This indicates a complete collapse of the population distribution, revealing that these methods achieve convergence primarily through aggressive exploitation, effectively sacrificing exploration. The Genetic Algorithm (GA) further illustrates this trade-off, exhibiting not only low diversity but also significantly worse (higher) free energy values across all transformations.
These findings strongly support the central hypothesis of this section: the dynamics of WE are fundamentally tied to the intrinsic geometry of the fitness landscape. The algorithm’s update rule—a balance between the objective function gradient ($\nabla f$) and the population density gradient ($\nabla\log\rho$)—enables it to adapt its search direction relative to the distribution of individuals, rather than relying on fixed, absolute coordinates. Consequently, when the search space is translated, scaled, or rotated, the relative relationship between the gradient forces and the entropic repulsion within the population is preserved. This allows WE to maintain a consistent balance between exploration and exploitation, leading to statistically invariant performance outcomes. This inherent invariance is a key advantage for real-world optimization problems where the optimal coordinate system or scaling is not known a priori, making WE a robust and versatile optimization framework.